find the solution of each system
y=7x+14
y=3x-6
also
4x+y=-11
2x-y=-1

Answer by darq(90)

y=7x+14
y=3x-6
then you can write

7x+14=3x-6
4x=-20
x=-5

y=3x-6
y=-15-6
y=-21

the second question

4x+y = -11
2x-y = -1
+____________
6x = -12
x=-2

if x is -2

2x-y=-1
-4-y=-1
y=-3